OpenStudy (anonymous):

Derive the equation of the parabola with a focus at (0, –4) and a directrix of y = 4. I understand a little bit of it, and... well, here is what I have so far... (x-h)^2 = 4p(y-k) h=0 k+p=-4 y=k-p=4 What's next?

OpenStudy (anonymous):

@triciaal

OpenStudy (triciaal):

from what you have h = 0 would mean x^2 = 4 p(y- k)

OpenStudy (triciaal):

give me a few minutes to get the diagram

OpenStudy (anonymous):

got it take your time

OpenStudy (triciaal):

" the equation of the parabola with focus (a , b) and directrix y = c is (x-a)^2 + b^2 -c^2 = 2(b-c)y"

OpenStudy (anonymous):

give it to me tomorrow, gtg

OpenStudy (triciaal):

Let ( x0 , y0 ) be any point on the parabola. Find the distance between (x0 , y0) and the focus. Then find the distance between (x0 , y0) and directrix. Equate these two distance equations and the simplified equation in x0 and y0 is equation of the parabola.

OpenStudy (triciaal):

distance between (x, y) and focus (0, -4) rt of (y + 4)^2 + x^2 distance between (x, y) and directrix = [y-4] set them equal and square each side